# SERA-H: Beyond Native Sentinel Spatial Limits for High-Resolution Canopy Height Mapping

### v2

@@ 2.4.2 Network Initialization and Optimization
The batch size was set to 2, with each training sample consisting of a sequence of 16 triplet images ($S2,S1_{asc},S1_{dsc}$), with each triplet image containing 14 channels. The ADAM optimizer (Kingma and Ba, 2017) was used with a learning rate of $1\times 10^{-4}$. The loss function employed was a Smooth L1 loss  [^6] https://docs.pytorch.org/docs/stable/generated/torch.nn.SmoothL1Loss.html. A ReduceLROnPlateau  [^7] https://docs.pytorch.org/docs/stable/generated/torch.optim.lr_scheduler.ReduceLROnPlateau.html scheduler was employed to reduce the learning rate by a factor of 0.1 whenever the validation loss failed to improve by at least 0.05 for 5 consecutive epochs. Early stopping was implemented with a patience of 10 and a minimum delta of 0.05 on the validation loss function. Training was performed on the Jean Zay supercomputer, using a node equipped with four NVIDIA Tesla V100 SXM2 32 GB GPUs and 40 Intel Cascade Lake 6248 CPUs.
@@ 2.5 Experimental Design
Table 1: Ablation study. Evaluation of the impact of super-resolution (e.g., EDSR compared to Bilinear Interpolation, BI), the use of time series (e.g., UTAE compared to U-Net), and the sequence length (4, 8, or 16 triplet images) on canopy height prediction performance in the test area of the dataset. Note that to limit computational overhead, all models presented here were trained on reduced input patches of $32\times 32$ pixels. Params denotes the total number of trainable parameters in millions. Best results are highlighted in bold.
Model Configuration Params MAE RMSE nMAE Tree Cover IoU
(M) (m) (m) (%) (%)
BI-UNet (mosaic) 31.0 3.39 4.90 27.55 83.57
BI-UTAE (16img) 1.1 3.35 4.84 26.78 83.91
@@
2.5.1 Ablation Study Setup
To evaluate the impact of different architectural components and identify the optimal configuration, we conducted an ablation study. The baseline architecture, denoted as SERA-H (EDSR-UTAE (16img)), combined the EDSR super-resolution module with the UTAE regression model using a sequence of 16 input images.
To assess the contribution of the super-resolution module, we trained a variant replacing EDSR with a standard $\times 4$ bilinear interpolation upsampling, denoted as BI-UTAE (16img). To analyze the influence of temporal modeling, two additional models employing a U-Net architecture instead of UTAE were evaluated. These models processed single composite images rather than time series and were designated as BI-UNet (bilinear interpolation upsampling) and EDSR-UNet (EDSR upsampling). These composites were generated using the temporal mean for Sentinel-1 and the temporal median for Sentinel-2. Finally, to determine the sensitivity to the sequence length, we trained two variants of SERA-H with reduced temporal inputs: eight images (EDSR-UTAE (08img)) and four images (EDSR-UTAE (04img)). To reduce computational overhead during this comparative analysis, all models in the ablation study were trained on smaller input patches of $32\times 32$ pixels.
2.5.2 Comparison with State-of-the-Art Methods
To evaluate the performance of SERA-H, we established a benchmarking protocol against five recent state-of-the-art (SOTA) methods: Fogel et al. (2025), Tolan et al. (2024), Liu et al. (2023), Schwartz et al. (2025), and Pauls et al. (2025). In the remainder of this paper, these models are referred to as Fogel, Tolan, Liu, Schwartz, and Pauls, respectively. These baselines differ in their supervision strategies and input data resolutions. On one hand, Schwartz and Pauls trained models using GEDI data on Sentinel-1 and Sentinel-2 imagery, producing height maps at a 10 m resolution. While Schwartz used standard temporal composites, Pauls exploited Sentinel-2 time series, using composites only for Sentinel-1. On the other hand, approaches based on commercial very high-resolution imagery all used data derived from ALS as a supervision reference: Tolan used 1 m Maxar imagery, Liu used 3 m PlanetScope imagery, and Fogel employed 1.5 m SPOT-6/7 images. Regarding the latter, it is worth noting that while SPOT imagery is generally commercial, the specific dataset used by Fogel is available free of charge for France through the Data Terra - DINAMIS  [^8] https://openspot-dinamis.data-terra.org/ institutional program.
To ensure a consistent quantitative evaluation at our reference resolution of 2.5 m, a standardization procedure was applied to all baseline predictions. Maps with a native resolution lower than the target (10 m and 3 m) were upsampled using bilinear interpolation, whereas maps with a finer native resolution (1.5 m and 1 m) were downsampled using max pooling to preserve the top canopy height information.
@@ 3.1 Ablation Study: Impact of Super-Resolution and Temporal Depth
These results demonstrate that the super-resolution module substantially improves prediction quality compared to standard bilinear interpolation (BI). This trend is first observed on architectures based on composite mosaics: the EDSR-UNet model outperformed BI-UNet, reducing the Mean Absolute Error (MAE) by $0.30$ m (from $3.39$ m to $3.09$ m) and the Root Mean Square Error (RMSE) by $0.34$ m. This performance gain was even more pronounced for temporal architectures: replacing bilinear interpolation with EDSR within the UTAE model resulted in a significant reduction in MAE of $0.62$ m (from $3.35$ m for BI-UTAE to $2.73$ m for SERA-H). These results confirm that training a dedicated upsampling module makes it possible to reconstruct the fine details of the canopy with much greater accuracy than conventional interpolation methods.
The use of the temporal dimension proved to be an equally important factor. Comparison of different architectures in Table˜1 showed that models with time series (UTAE) consistently outperformed those based on static annual composites (U-Net). This efficiency was evident even in the configuration without super-resolution: the BI-UTAE model performed better than BI-UNet (MAE of $3.35$ m vs. $3.39$ m), while being notably lighter in terms of the number of parameters ($1.1$ million vs. $31$ million). The gap widened further with the addition of super-resolution: the SERA-H temporal model improved the MAE by $0.36$ m compared to its static equivalent EDSR-UNet ($2.73$ m vs. $3.09$ m). These results confirm that the use of time series is a more relevant and effective approach than the use of mosaics for canopy height mapping.
Finally, the ablation study revealed a direct positive correlation between the length of the input time series and the accuracy of the model. Reducing the number of images led to a gradual deterioration in performance. Specifically, the MAE increased by $0.22$ m when using 8 instead of 16 images (from $2.73$ m to $2.95$ m), and worsened by an additional $0.09$ m when retaining only 4 images ($3.04$ m). This confirmed that the model benefited from dense temporal sampling to resolve ambiguities in canopy structure. Given that the average availability per area in our dataset was $13.51$ triplet images, the configuration of 16 triplets already exceeded this average. We therefore considered that exploring a model requiring an even longer sequence (e.g., 32 triplets) was not relevant.
@@ 3.2 Comparison with State-of-the-Art (SOTA)
Model Input Reference MAE RMSE nMAE Tree Cover IoU
Images Images (m) (m) (%) (%)
@@ 3.2.1 Qualitative Evaluation
Despite utilizing the same coarse-resolution Sentinel data, SERA-H generated predictions that surpassed the native resolution of the input. It produced clearer and more precise contours of forests and individual trees, allowing us to approach the scale of single-tree detection, which is significantly superior to other Sentinel-based products. This enhancement was largely driven by the use of ALS data as a reference, which provided a much higher frequency signal, enabling sharper and more accurate predictions. This confirms the utility of upsampling Sentinel images via a dedicated super-resolution method to align with the fine granularity of ALS data.
@@
Liu trained their model on a much broader dataset covering all of Europe; consequently, their model is not specifically tailored to the French forest types targeted in our test dataset. Despite this, Liu’s model relies on commercial PlanetScope imagery, which has a resolution more than three times finer ($3$ m) than that of Sentinel. However, even against this higher-resolution input, SERA-H produced sharper, more precise predictions with clearer delineation of forest boundaries and individual trees.
The comparison with Fogel is the most relevant, as it is based on the same training dataset (Open-Canopy) and the same ALS-based canopy height references. Although Fogel employed SPOT-6/7 imagery, which offers a spatial resolution nearly seven times higher ($1.5$ m) than our Sentinel inputs, SERA-H achieved a comparable prediction performance. While this resolution gap logically grants Fogel’s maps a finer level of textural detail, the ability of SERA-H to approach these performance metrics demonstrates the effectiveness of our super-resolution approach in bridging the gap between open-access and commercial imagery.
@@ 3.2.2 Quantitative Evaluation
In this section, we present the quantitative evaluation of SERA-H alongside other state-of-the-art models. Table˜2 summarizes the performance metrics (detailed in Section˜2.5.3) on the test dataset. To provide a more detailed analysis of the prediction errors, Fig.˜5 illustrates the distribution of residuals across height bins, while Fig.˜6 displays density scatterplots against the ALS reference data.
@@
As noted in Section˜3.2.1, comparing our results with Pauls and Schwartz should be done with some caution, given their 10 m native resolution and reliance on GEDI reference. Metric discrepancies are therefore to be expected due to these differences in resolution and reference. However, as the height maps derived from ALS constitute the most reliable local estimate of canopy height at a very high resolution, it remains relevant to evaluate all models against this reference. Furthermore, this comparison allows us to identify the methodological benefit of SERA-H, revealing the accuracy improvement over existing methods that utilize Sentinel imagery as input.
The results reported in Table˜2 are consistent with the improved spatial detail observed in the maps (Section˜3.2.1). SERA-H showed significantly lower errors than the Sentinel- and GEDI-based references: we obtained an MAE of $2.60$ m, compared to $4.47$ m for Schwartz and $5.13$ m for Pauls. The relative error (nMAE) was also reduced by half ($20.40\%$ compared to over $42\%$ for Pauls). Furthermore, with a Tree Cover IoU of $87.25\%$ compared to a range of $35$-$45\%$ for the other models, SERA-H demonstrated a much greater ability to correctly delineate vegetation. Finally, we verified these trends by calculating the metrics at a resolution of 10 m (A), corresponding to the native resolution of Sentinel images and related products. Although performance differences were reduced with spatial aggregation, the hierarchy of results remained the same. For example, for MAE, SERA-H obtained $3.00$ m, while Pauls obtained $3.36$ m and Schwartz obtained $3.72$ m, confirming the robustness of our model’s predictions.
Further intercomparison analysis (Fig.˜5 and Fig.˜6) confirmed the good performance of the SERA-H model results. It highlights that models based on the Sentinel/GEDI pair systematically showed significantly greater error dispersion, as illustrated by more diffuse scatter plots and wider interquartile ranges. In addition, a strong tendency toward overestimation could be observed across most of the spectrum (0-$30$ m), where the median errors were above zero. While performance became broadly comparable for the 2020-$30$ m height class and the model by Pauls showed a lower median bias for trees exceeding 30 m, SERA-H nevertheless confirmed its overall superiority over models derived from Sentinel images.
@@
Quantitative analysis (Table˜2) shows that SERA-H significantly outperforms the approaches of Tolan and Liu. Despite the theoretical advantage conferred by the use of very high-resolution commercial images (PlanetScope at 3 m resolution and Maxar at 1 m), these models yielded Mean Absolute Errors (MAE) of $4.40$ m and $5.49$ m, respectively, which were higher than the $2.60$ m obtained by SERA-H. This hierarchy is visually confirmed by the scatter plots (Fig.˜6): where the predictions of Liu and Tolan showed high dispersion and struggled to follow the 1:1 line (with $R^{2}$ values of $0.581$ and $0.376$), SERA-H showed strong linearity ($R^{2}=0.818$). While this discrepancy could be explained partly by domain shift, given that Tolan was trained in the United States and Liu across Europe, these results suggest that excessive geographical generalization can be detrimental to local accuracy.

### v3

@@ 2.4.2 Network Initialization and Optimization
The batch size was set to 2, with each training sample consisting of a sequence of 16 triplet images $(S2,S1_{\text{asc}},S1_{\text{dsc}})$, with each triplet image containing 14 channels. The ADAM optimizer (Kingma and Ba, 2017) was used with a learning rate of $1\times 10^{-4}$. The loss function employed was a Smooth L1 loss  [^8] https://docs.pytorch.org/docs/stable/generated/torch.nn.SmoothL1Loss.html. A ReduceLROnPlateau  [^9] https://docs.pytorch.org/docs/stable/generated/torch.optim.lr_scheduler.ReduceLROnPlateau.html scheduler was employed to reduce the learning rate by a factor of 0.1 whenever the validation loss failed to improve by at least 0.05 for 5 consecutive epochs. Early stopping was implemented to halt training if the validation loss failed to improve by at least 0.05 for 10 consecutive epochs. Training was performed on the Jean Zay supercomputer, using a node equipped with four NVIDIA Tesla V100 SXM2 32 GB GPUs and 40 Intel Cascade Lake 6248 CPUs.
@@ 2.5 Experimental Design
2.5.1 Ablation Study Setup
To evaluate the impact of different architectural components and identify the optimal configuration, we conducted an ablation study. Our proposed architecture, denoted as SERA-H (EDSR-UTAE (16img)), consists of the EDSR super-resolution module integrated with the UTAE regression model, utilizing a sequence of 16 input images. To reduce computational overhead during this comparative analysis, all models in the ablation study were trained on smaller input patches of $32\times 32$ pixels.
To assess the contribution of the super-resolution module, we trained a variant replacing EDSR with a standard $\times 4$ bilinear interpolation upsampling, denoted as BI-UTAE (16img). To ensure a fair comparison, the capacity of this variant was scaled to match the number of parameters of the full model by increasing the width of the UTAE encoder and decoder layers and expanding the attention mechanism dimensions. This guarantees that any performance gains are attributable to the architectural design rather than a simple increase in total capacity. This increase in the number of model parameters is indicated by the asterisk appended to the model name.
To analyze the influence of temporal modeling, two additional models employing a U-Net architecture instead of UTAE were evaluated: BI-UNet (bilinear interpolation upsampling) and EDSR-UNet (super-resolution upsampling). These models processed single composite images instead of full time series. These composites were generated by calculating the temporal mean for Sentinel-1 and the temporal median for Sentinel-2. Similarly, the number of parameters in the BI-UNet model was increased to match that of EDSR-UNet to ensure a fair architectural evaluation.
To demonstrate the advantage of supervised learning at the target 2.5 m resolution instead of the native 10 m resolution, we trained a variant denoted as EDSR-UTAE-10m (16img). This model follows exactly the same architecture as SERA-H, except for the upsampling layer in the super-resolution module, which has been removed, yielding outputs at 10 m resolution. During training, the ALS reference resolution was downsampled to 10 m using max-pooling to match the model’s output. For evaluation, the 10 m predictions were subsequently upsampled to 2.5 m using bilinear interpolation.
Finally, to determine sensitivity to sequence length, we trained two variants of SERA-H with reduced temporal length inputs: an eight-image model (EDSR-UTAE (08img)) and a four-image model (EDSR-UTAE (04img)).
2.5.2 Comparison with State-of-the-Art Methods
To evaluate the performance of SERA-H, we established a benchmarking protocol against five recent state-of-the-art (SOTA) methods: Fogel et al. (2025), Tolan et al. (2024), Liu et al. (2023), Schwartz et al. (2025), and Pauls et al. (2025). In the remainder of this paper, these models are referred to as Fogel, Tolan, Liu, Schwartz, and Pauls, respectively. These baselines differ in their supervision strategies and input data resolutions. On one hand, Schwartz and Pauls trained models using GEDI data on Sentinel-1 and Sentinel-2 imagery, producing height maps at a 10 m resolution. While Schwartz used standard temporal composites, Pauls exploited Sentinel-2 time series, using composites only for Sentinel-1. On the other hand, approaches based on commercial very high-resolution imagery all used data derived from ALS as a supervision reference. For input data, Tolan used 1 m Maxar imagery, Liu utilized 3 m PlanetScope imagery, and Fogel employed 1.5 m SPOT-6/7 images. Regarding the latter, it is worth noting that while SPOT imagery is generally commercial, the specific dataset of satellite images used by Fogel is available free of charge for France through the Data Terra - DINAMIS  [^10] https://openspot-dinamis.data-terra.org/ institutional program. Furthermore, it is important to specify that Fogel was trained on the Open-Canopy benchmark as reference data, whereas the other ALS-supervised models relied on their own specific LiDAR datasets: Tolan focused on the USA, while Liu utilized a generalized dataset across Europe.
To ensure a consistent quantitative evaluation at our reference resolution of 2.5 m, a standardization procedure was applied to all baseline predictions. Maps with a native resolution lower than the target (Pauls, Schwartz 10 m and Liu 3 m) were upsampled using bilinear interpolation, whereas maps with a finer native resolution (Fogel 1.5 m and Tolan 1 m) were downsampled using max pooling to preserve the top canopy height information.
@@ 2.5.2 Comparison with State-of-the-Art Methods
Table 1: Ablation study. Evaluation of the impact of super-resolution (EDSR vs. Bilinear Interpolation, BI), temporal modeling (UTAE vs. U-Net), sequence length (4, 8, or 16 triplets), and supervision resolution (2.5 m vs. Sentinel native 10 m) on canopy height prediction. All metrics are evaluated in the test area at a standardized resolution of 2.5 m (model with a 10-meter output was oversampled for comparison). Note that to limit computational overhead, all models presented here were trained on reduced input patches of $32\times 32$ pixels. Params denotes the total number of trainable parameters in millions. Best results are highlighted in bold.
Model Configuration Params MAE RMSE nMAE Tree Cover
(M) (m) (m) (%) IoU (%)
BI-UNet (mosaic) 75.8 3.54 5.10 28.87 83.09
BI-UTAE (16img) 49.9 3.12 4.58 25.20 84.85
@@ 3.1 Ablation Study: Impact of Super-Resolution and Temporal Depth
These results demonstrate that the super-resolution module substantially improves prediction quality compared to standard Bilinear Interpolation (BI). This trend is first observed in architectures based on composite mosaics: the EDSR-UNet model outperformed BI-UNet, reducing the Mean Absolute Error (MAE) by $0.45$ m (from $3.54$ m to $3.09$ m) and the Root Mean Square Error (RMSE) by $0.54$ m. This performance gain is confirmed for temporal architectures: replacing BI with EDSR within the UTAE model resulted in a significant MAE reduction of $0.39$ m (from $3.12$ m for BI-UTAE to $2.73$ m for SERA-H). These findings confirm that a dedicated upsampling module enables the reconstruction of fine canopy details with significantly greater accuracy than conventional interpolation methods.
The use of the temporal dimension proved to be an equally important factor. Comparison of different architectures in Table˜1 showed that models with time series (UTAE) consistently outperformed those based on static annual composites (U-Net). This efficiency was evident even in the configuration without super-resolution: the BI-UTAE model performed better than BI-UNet (MAE of $3.12$ m vs. $3.54$ m), while being notably lighter in terms of the number of parameters ($49.9$ million vs. $75.8$ million). The difference between these two methods is also visible on models that perform super-resolution: the SERA-H temporal model improved the MAE by $0.36$ m compared to its static equivalent EDSR-UNet ($2.73$ m vs. $3.09$ m). These results demonstrate that exploiting temporal redundancy does indeed allow for better characterization of forest structure than the use of static mosaics.
The significant deterioration in scores for the EDSR-UTAE-10m variant, which has the highest MAE in the benchmark at 3.89 m, demonstrates that degrading the resolution of ALS references leads to a major loss of structural details, confirming the value of high-resolution training targets. This performance gap also proves that SERA-H performs true spatial reconstruction and is not simply a disguised version of a 10 m prediction followed by bilinear interpolation.
Finally, the ablation study revealed a direct positive correlation between the length of the input time series and the accuracy of the model. Reducing the number of images led to a gradual deterioration in performance. Specifically, the MAE increased by $0.22$ m when using 8 images (EDSR-UTAE (8img)) instead of 16 with SERA-H (from $2.73$ m to $2.95$ m), and worsened by an additional $0.09$ m when retaining only 4 images ($3.04$ m with EDSR-UTAE (4img)). This confirmed that the model benefited from dense temporal sampling to resolve ambiguities in canopy structure. Given that the average availability per area in our dataset was $13.51$ triplet images, the configuration of 16 triplets already exceeded this average. We therefore considered that exploring a model requiring an even longer sequence (e.g., 32 triplets) was not relevant.
@@ 3.2.1 Qualitative Evaluation
Model Input Reference MAE RMSE nMAE Tree Cover
Images CHM (m) (m) (%) IoU (%)
@@
Despite utilizing the same coarse-resolution Sentinel data, SERA-H generated predictions that surpassed the native resolution of the input. It produced clearer and more precise contours of forests and individual trees, allowing us to approach the scale of the individual tree, which is significantly superior to other Sentinel-based products. This enhancement was largely driven by the use of ALS data as a reference, which provided a much higher frequency signal, enabling sharper and more accurate predictions. This confirms the utility of upsampling Sentinel images via a dedicated super-resolution method to align with the fine granularity of ALS data.
@@
Liu trained their model on a much broader dataset covering all of Europe; consequently, their model is not specifically tailored to the French forest types targeted in our test dataset. Despite this, Liu’s model relies on commercial PlanetScope imagery, which has a resolution more than three times finer ($3$ m) than that of Sentinel. However, even against this higher-resolution input, SERA-H equals or even surpasses Liu’s spatial definition (definition of forest and isolated trees). It can also be observed that, in many areas, SERA-H more accurately reflects internal variations in forest structure than Liu’s predictions (variations in tree height and gaps in the forest).
The comparison with Fogel is the most relevant, as it is based on the same training dataset and ALS-based canopy height references (Open-Canopy dataset). Although Fogel employed SPOT-6/7 imagery, which offers a spatial resolution nearly seven times higher ($1.5$ m) than our Sentinel inputs, SERA-H achieved very close prediction performance. While this resolution gap logically grants Fogel’s maps a finer level of textural detail, the ability of SERA-H to approach these performance metrics demonstrates the effectiveness of our super-resolution approach in bridging the gap between open-access and commercial imagery.
@@ 3.2.2 Quantitative Evaluation
In this section, we present the quantitative evaluation of SERA-H alongside other state-of-the-art models. Table˜2 summarizes the performance metrics (detailed in Section˜2.5.3) on the test dataset. To provide a more detailed analysis of the prediction errors, Fig.˜5 illustrates the distribution of residuals across height bins, while Fig.˜6 displays density scatterplots against the ALS reference data of the test dataset.
@@
As noted in qualitative results (Section˜3.2.1), comparing our results with Pauls and Schwartz should be done with some caution, given their 10 m native resolution and reliance on GEDI reference. Metric discrepancies are therefore to be expected due to these differences in resolution and reference. However, as the height maps derived from ALS constitute the most reliable local estimate of canopy height at a very high resolution, it remains relevant to evaluate all models against this reference. Furthermore, this comparison allows us to identify the methodological benefit of SERA-H, revealing the accuracy improvement over existing methods that utilize Sentinel imagery as input.
The results reported in Table˜2 are consistent with the improved spatial detail observed in the qualitative results (Section˜3.2.1). SERA-H showed significantly lower errors than the Sentinel- and GEDI-based references: we obtained an MAE of $2.60$ m, compared to $4.47$ m for Schwartz and $5.13$ m for Pauls. The relative error (nMAE) was also reduced by half ($20.40\%$ compared to over $42\%$ for Pauls). Furthermore, with a Tree Cover IoU of $87.25\%$ compared to a range of $35$-$45\%$ for the other models, SERA-H demonstrated a much greater ability to correctly delineate vegetation. Finally, we verified these trends by calculating the metrics at a resolution of 10 m (A), corresponding to the native resolution of Sentinel images and related products. Although performance differences were reduced with spatial aggregation, the hierarchy of results remained the same. For example, for MAE, SERA-H obtained $3.00$ m, while Pauls obtained $3.36$ m and Schwartz obtained $3.72$ m, confirming the robustness of our model’s predictions.
Further intercomparison analysis (Fig.˜5 and Fig.˜6) confirmed the good performance of the SERA-H model results. It highlights that models based on the Sentinel/GEDI pair systematically showed significantly greater error dispersion, as illustrated by more diffuse scatter plots and wider interquartile ranges. In addition, these sentinel/GEDI models show a strong tendency toward overestimation across most of the spectrum (0-$30$ m), where the median errors were above zero. However, their performance becomes broadly comparable to that of SERA-H within the 20 to 30 m class, and Pauls’ model demonstrated a lower median bias for trees exceeding 30 m. While all models eventually encounter signal saturation as canopy height increases, the architecture proposed by Pauls appears to be the most resilient to this phenomenon. This greater resilience to saturation allows it to maintain more reliable predictions for tall trees, ultimately outperforming other models in this specific range. Nevertheless, SERA-H confirmed its overall superiority over models derived from Sentinel images across the rest of the height range.
@@
Quantitative analysis (Table˜2) shows that SERA-H significantly outperforms the approaches of Tolan and Liu. Despite the theoretical advantage conferred by the use of very high-resolution commercial images (PlanetScope at 3 m resolution and Maxar at 1 m), these models yielded Mean Absolute Errors (MAE) of $4.40$ m and $5.49$ m, respectively, which were higher than the $2.60$ m obtained by SERA-H. This hierarchy is visually confirmed by the scatter plots (Fig.˜6): where the predictions of Liu and Tolan showed high dispersion and struggled to follow the 1:1 line (with $R^{2}$ values of $0.581$ and $0.376$), SERA-H showed strong linearity ($R^{2}=0.818$). While this discrepancy could be explained partly by domain shift, given that Tolan was trained in the United States and Liu across Europe, these results suggest that excessive geographical generalization can be detrimental to local accuracy (Fischer et al., 2025).